OpenStudy (anonymous):

four inches of heavy wet snow are equilvalient to 2 inches of rain. Estimate the water content in 15 inches of heavy wet snow. someone was helping me and I lost them or something

OpenStudy (anonymous):

I think this is just comparing ratios: 4/2 = 15/x, solve for x.

OpenStudy (anonymous):

4in (snow) = 2in (rain) 15in (snow) = 2in(rain)/4in(snow) =15/2 in (rain) = 7.5 in of rain

OpenStudy (anonymous):

sorry, 15in * 2/4

OpenStudy (anonymous):

Okay so what we know is this: 4 inches of snow = 2 inches of rain. So let's find out what the ratio between heavy snow and rain. We are finding how much rain is in snow, so let's put that at at the top. 1.) 2 inches of rain / 4 inches of snow = 0.5 rain/snow Okay, so now we have a ratio to work with. How much snow did it say we had? 15 inches. So, let's multiple this by our ratio. 2.) 15 inches of snow * 0.5 rain/snow = 7.5 inches of rain. *Note* *Make sure you keep track of those units. In this case, the units are inches of snow and inches of rain. When we took out ratio, "inches of" canceled each other out. However, rain/snow did not cancel each other out, because they aren't the same units. As you see, in step 2 rain/snow was carried down. This is an important lesson to learn, because it will always come back to you in math through seconds/minutes/hours, inches/feet/yards/miles, mm/cm/m/km, etc. So in step 2, "inches of" did not cancel out, but snow did. So we were left with "inches of rain" as our final unit.* 7.5 inches of rain is your answer. I hope that helps.